# leftmargin=0.15cm, rightmargin=0.15cm, font=italic, vskip=0.10cm, indentfirst=false

### v4

@@ IV-C Instrument Design & Development
    Used a snowball review of top security–conference works on signing primitives [11, 71] and empirical studies (see section II), to identify gaps and avoid redundant questions.
Next, we refined our interview protocols through a series of practice and pilot interviews[72]. We first conducted two practice interviews with the secondary authors of this work, followed by pilot interviews with the first two interview subjects. For example, we moved the question about each participant’s team’s major product to the demographics section to provide context for their responses. Our full interview instrument consisted of three main topics (B-D). Topics B and C did not fit into the same analysis framework as Topic D. Thus, following the guidance of Voils et al. [73], we analyzed those topics in another work [BLINDED]. [^1] That work is blinded for review. We attest no duplication of material. In this work, we focus on the data for Topic D.
A summary of the questions in our interview protocol is shown in Table I. Topic D consisted of five main question categories. Questions D3–D4 probe current tooling experiences. Questions D1–D2 explore previous tools, and D5 asks about transition factors. See section -B for the full protocol.
TABLE I: Summary of interview protocol. Data from Topics B and C are analyzed in [BLINDED]. We analyze Topic D in this paper. We present our full protocol in section -B
Topic(# Questions) Sample Questions
A. Demographics (4) What is your role in your team?
@@
B. Perceived software supply chain risks (4) Can you describe any specific software supply chain attacks (e.g., , incidents with 3rd-party dependencies, code contributors, OSS) your team has encountered?
C. Mitigating risks with signing (8) How does the team use software signing to protect its source code (what parts of the process is signing required)?
D. Signing tool adoption (5)
Question Categories in Topic D
@@ D. Signing tool adoption (5)
D1. What factors did the team consider before adopting this tool/method (Sigstore) over others?
@@
D2. What was the team’s previous signing practice/tool before the introduction of Sigstore?
@@
D3. How does your team implement Sigstore (which components of Sigstore does the team mostly use)?
D4. Have you encountered any challenge(s) using this tool of choice?
@@
D5. Have you/your team considered switching Sigstore for another tool?
IV-D Data Collection
Population: Our study investigates signing tools using a case study of the Sigstore ecosystem. As such, our target participant pool is centered on Sigstore users. Furthermore, our research questions, as reflected in our interview instrument, aim to understand the reasons prior to adoption that teams choose software tools (Sigstore), how Sigstore is used in signing, and the challenges and benefits of Sigstore.
To effectively answer these questions related to decision-making, we required a target population of high-ranking or experienced practitioners who either oversee Software signing compliance or manage the infrastructure of their respective teams or organizations.
@@ IV-D Data Collection
To recruit this target population, we employed a non-probability purposive and snowball sampling approach  [74]. We leveraged the personal network of one of the authors to send an initial invitation to members of the 2023 KubeCon (hosted by the Linux Foundation) organizing committee, who were also part of the In-toto [75, 76] Steering Committee (ITSC). This initial invitation yielded 6 participants. We then expanded our recruitment through snowball sampling, relying on recommendations from initial participants (5) and authors’ contacts (7) to recruit an additional 12 participants. We interviewed 18 subjects, but only 17 participants could share enough details for analysis.
@@
Participant Demographics: Our subjects were experienced security practitioners responsible for initiating or implementing their organization’s security controls or compliance. This gives them the relevant context to assess their organization’s strategies for adopting software signing tools. Subjects came from 13 distinct organizations, all companies. Relevant demographics are in Table II. Thirteen participants reported their organizations use Sigstore. Two use internally developed tools and two rely on Notary v1 and PGP signing. Those four participants enrich our findings, as their experiences shed light on usability factors that hinder adoption.
@@
TABLE II: Demographics. For Role, “Senior management” are senior managers/directors/executives; “Technical leader” are senior/lead/partner/principal engineers; and “Engineer” and “Manager” are junior staff.
@@
S13 Senior mgmt. 16 years Internal security tool
@@
[…]
IV-E Data Analysis
We transcribed our interview recordings using www.rev.com’s human transcription service and anonymized all potentially identifiable information.
After transcription and anonymization, our analysis began with a data familiarization stage [77] and proceeded through four phases, carried out by two analysts. [^2] One of the analysts is the lead author. In the first two phases, we inductively memoed and coded our transcripts, developing our initial codebooks without referencing our usability framework [78]. In the final phase, we derived themes from our inductively generated codes and then applied our usability framework to further refine the extracted themes. We describe these steps in detail next. We remind the reader that the transcript memoing and codebook was developed across the full interview transcripts, including Topics B and C. This manuscript focuses on the data and codes pertinent to Topic D.
IV-E1 Memoing & Codebook Creation
@@
First, beginning with the initial 18 transcripts, the two analysts independently memoed a subset of six anonymized transcripts across three rounds of review and discussion, following the recommendations and techniques outlined by O’Connor & Joffe [79] and Campbell et al. [80]. This approach ensured the reliability of our analysis process while optimizing data resources. [^3] Although there is no consensus on the ideal proportion of data to use, O’Connor and Joffe suggest that selecting 10–25% is typical [79]. We randomly chose 6 transcripts (33% of data) to develop our initial codebook. After each round, emerging codes were discussed in meetings, resulting in a total of 506 coded memos. We then reviewed the coded memos and finalized an initial codebook with 36 code categories.
@@
To assess the reliability of the coding process, we combined the techniques outlined by Maxam & Davis [81] and Campbell et al. [80]. Specifically, we randomly selected an unanalyzed transcript, which we coded independently. Following Feng et al.’s [82] recommendation, we used the percentage agreement measure to evaluate consistency, as both analysts had contributed to developing the codebook. We obtained an 89% agreement score, suggesting a stable coding scheme [80, 79]. We resolved disagreements through discussion.
IV-E2 Codebook Revision
@@
Following the reliability assessment of our codebook, and given the substantive agreement recorded, one of the analysts proceeded to code the remaining transcripts in two batches of six transcripts per round. In the first batch, six new code categories were introduced, which were subsequently reviewed by the other analyst. The same analytical process was applied to the final six transcripts, leading to the addition of four new code categories, which were subsequently reviewed by the other analyst. Notably, of the 10 codes added during this phase of codebook refinement, only three provided new insights into the data. By mutual agreement between both analysts, these were categorized as minor codes.
@@
To assess the adequacy of our sample size and analysis, we calculated code saturation for each category in our codebook. Following the recommendations of Guest et al. [83], we measured the cumulative number of new codes introduced in each interview. We saw saturation after interview 12, with no new codes after that. The results are presented in Figure 7.
IV-E3 Thematic Analysis & Usability Framework Analysis
@@
Each interview had a median of 15 codes for topic D. Next, following the method proposed by Braun & Clarke [84], we derived initial themes from our code categories. We focused on identifying themes related to Topic D of our protocol (Table I).
We then conducted another round of analysis using a formative usability framework. We selected Cresswell’s framework (Figure 1) based on the nature of our data.
V Results
@@
First, we present a Sankey diagram illustrating the changes in signing tool usage among our participants’ teams (Figure 4). Many participants (6) transitioned from using no tools to adopting Sigstore. Others transitioned from existing tools to Sigstore: 3 from GPG, 2 from Notary, 1 from Skopeo, and 2 from proprietary or internal tools. The final 4 subjects did not transition to Sigstore.
@@
To understand why they did (and did not) change tools, we structure the results by RQ using Cresswell’s framework.
V-A RQ1: Strengths & Weaknesses of Sigstore
@@
V-A1 Sigstore’s Strength
We present a summarized list of practitioner-reported advantages of using Sigstore in Table III. First, we extracted the benefits our participants experienced while using Sigstore from their responses. These reflections capture their firsthand experiences with Sigstore.
TABLE III: Summary of Practitioner-Reported Advantages of Using Sigstore. Practitioners cited Sigstore’s technological and macroenvironmental factors as advantages.
@@ V-A1 Sigstore’s Strength
Ease of Use 8
Signing Workflow & Verification P1, P2, P7, P9, P14-17
Setting up with automated CI/CD actions P9
No key distribution problems P1
Use of Short-lived Keys & Certificate 3 –P2, P3, P15
Signer ID Management 4
Use of OIDC(Keyless) to authenticate signers P3, P5, P10, P12
Compatibility with Several New Technologies 4
Integrability with SLSA build P16
Integrability with several container registries/technologies P12, P14
Integrability with several cloud-native applications P15
Precence of a Transparency Log 3
Transparency logs increase security P5, P14
Evaluation of signing adoption using logs P9
Bundling Signatures With Provenance Attestations 2 – P3, P4
Reliability of Service 1 – P7
@@
Free/Open-Source 2 – P7, P17
We grouped their responses into eight categories: ease of use, identity management for signers, compatibility with emerging technologies, use of short-lived keys and certificates, implementation of a transparency log, encouragement of bundling signatures with attestations, free access to the public Sigstore instance, and reliability of the Sigstore service. When mapped to the Cresswell framework, these factors fell under two categories: Technological and Macroenvironmental.
@@
Ease of use: This was the most frequently discussed advantage of Sigstore over other Software signing tools. Participants commonly highlighted the simplicity of following the Sigstore workflow—particularly how key management is fully automated reduce configuration overhead. P7 (security tools, cloud security) states, “ I don’t really know of any other options that provide the same conveniences that Sigstore provides… It really is just one command to sign something and then one more command to verify it, and so much of the work is handled by Sigstore.” Other advantages associated with Ease of Use, as mentioned by practitioners, include the seamless integration of the Sigstore workflow into CI/CD automation using GitHub Actions and the reduced complexity of key management.
@@
Compatibility with Several New Technologies: A portion of our participants reported that Sigstore’s integration and compatibility with a wide range of modern technologies provided a significant advantage. The mentioned technologies included security tooling such as SLSA build attestations, as well as containerization technologies and cloud computing applications. In most cases, the resources required to set up connections to these technologies were minimal or nonexistent. P12 (consultancy, cloud OSS security) states, “a lot of clients are getting into signing their containers pretty trivial to do if you use something like cosign…Sigstore is the easy choice because it just works on any registry, and integrating it to Kubernetes is pretty easy ”
Use of Short-lived Keys/Certificates: Sigstore’s enforcement of short-lived keys and certificates as a security feature was recognized by our participants as a major strength of the tooling. Traditional software signing tools require users to manually set expiration times for cryptographic elements such as signatures, keys, and certificates at the time of generation. While this approach provides customization flexibility, it also introduces the risk of long-lived cryptographic materials remaining valid for extended periods, increasing the likelihood of compromise. Participants appreciated Sigstore’s automated management of key and certificate lifetimes, which mitigates this risk. P2 (SAAS, SSC security tool) states, “So (with) PGP, you still have to figure out the key distribution problem. Because Sigstore uses short-lived keys that solves a lot of those issues around key storage and key distribution…”
Presence of a Transparency Log: The ability to audit signing actions was crucial for adopters. Participants valued the transparency log in the Sigstore ecosystem because it provides a tamper-resistant record of all changes to an artifact’s metadata—functionality not offered by other signing systems. This auditability gave users greater confidence in the integrity of signed artifacts. {quoting} P9 (SAAS, SSC security): “From a higher level, we’re going to use a transparency log to determine what was actually signed and what was written into that transparency log.”
Other technological factors noted by participants are the reliability of Sigstore’s public instance.
@@
Free/Open-Source: Sigstore’s open-source nature was also highlighted by participants as a key advantage. They noted that this reduces both setup and maintenance barriers, particularly for users leveraging Sigstore’s public deployments. {quoting} P7 (security tools, cloud security): “I think it’s amazing that Sigstore is free and public”
V-A2 Sigstore’s Weaknessses
We summarize the difficulties reported by participants in using Sigstore in Table IV. Most of these weaknesses mapped to multiple usability factors.
TABLE IV: Reported Difficulties Encountered by Participants in Using Sigstore. We indicate the associated usability factor of each category of weakness using – T-technology, P-social/human, O-organizational, M-macroenvironmental factors.

### v5

@@ 4.3 Instrument Design & Development
    Used a snowball review of top papers on signing primitives [104, 55] and relevant empirical studies (see section˜2) to identify gaps and avoid redundant questions. We seeded the review with works from prominent cybersecurity and software engineering venues (e.g., USENIX Security, IEEE S&P, ICSE, FSE) published in the past 10 years.
Next, we refined our interview protocols through a series of initial interviews [20]. We first conducted practice interviews with the secondary authors of this work, followed by pilot interviews with the first two interview subjects.
Table˜1 summarizes our final interview protocol. After the demographic section (Topic A), our interview instrument consisted of two main topics (B & C). Topic B comprised contextual questions about the supply chain risks that prompt signing and how signing is implemented across the software production process (e.g., where signing is required vs. optional, which artifacts are signed, and relevant workflow/policy constraints). Topic C presented questions to assess the usability of the signing tool(s) used by the participant’s organization, including perceived strengths and weaknesses, adoption rationale, implementation details, challenges, and any considerations about switching tools.
Table 1: Summary of interview protocol. We present our full protocol in Appendix˜B
Topic Sample Questions
A. Demographics What is your role in your team?
@@
B. Contextual Information (Risks & Use of Signing) Can you describe any specific software supply chain attacks (e.g., incidents with 3rd-party dependencies, code contributors, OSS) your team has encountered? How does the team use software signing to protect its source code (what parts of the process is signing required)?
C. Signing Tool Usability Questions
@@
C1. What factors did the team consider before adopting this tool/method (Sigstore) over others?
@@
C2. What was the team’s previous signing practice/tool before the introduction of Sigstore?
@@
C3. How does your team implement Sigstore (which components of Sigstore does the team mostly use)?
C4. Have you encountered any challenge(s) using this tool of choice?
@@
C5. Have you/your team considered switching Sigstore for another tool?
4.4 Data Collection
Population: Our study investigates signing tools using a case study of the Sigstore tool ecosystem. Accordingly, our participant pool focused on current and prospective users of this tool. Furthermore, since our research questions pertain to organizational decision-making, we targeted expert practitioners who either oversee Software signing compliance or manage the infrastructure of their respective teams or organizations.
@@ 4.4 Data Collection
To recruit this target population, we employed a non-probability purposive and snowball sampling approach  [9]. We leveraged the personal network of one of the authors to send an initial invitation to members of the 2023 KubeCon (hosted by the Linux Foundation) organizing committee, who were also part of the In-toto [94, 38] Steering Committee (ITSC). This initial invitation yielded 6 participants. We then expanded our recruitment through snowball sampling, relying on recommendations from initial participants (4) and authors’ contacts (7) to recruit an additional 11 participants. We report on data from 17 participants. [^1] We conducted 18 interviews. Participant 18 was a senior security expert. While aware of the organization’s overall use of signing, they reported limited knowledge of the signing tool’s specific operation and deployment.
@@
Participant Demographics: Our subjects were experienced security practitioners responsible for initiating or implementing their organization’s security controls or compliance. This gives them the relevant context to assess their organization’s strategies for adopting software signing tools. Subjects came from 13 distinct organizations, all companies. Relevant demographics are in Table˜2. Thirteen participants reported their organizations use Sigstore. Two use internally developed tools and two rely on Notary v1 and PGP signing. Those four participants enrich our findings, as their experiences shed light on usability factors that hinder adoption of sigstore by non-users.
@@
Table 2: Demographics. For Role, “Senior management” are senior managers/directors/executives; “Technical leader” are senior/lead/partner/principal engineers; and “Engineer” and “Manager” are junior staff. Table˜10 gives organizational info.
@@
P13 Senior mgmt. 16 years Internal security tool
@@
Interviews: We conducted our interviews over Zoom. Each interview lasted $\sim$50 minutes. The lead author conducted these interviews. We offered each subject a \$100 gift card as an incentive in recognition of their expertise.
Survey Attempt: We recognize the importance of data triangulation in case studies and therefore fielded a follow-on survey to complement our interviews. We deployed a survey to validate and model the prevalence and stability of our interview themes over time and across contexts. Despite five months of repeated solicitation in the Sigstore community, we received only 13 completed responses, of which 7 were sufficiently complete for analysis (see Appendix˜E). This regrettably echoes the well-documented difficulty of recruiting participants from highly specialized populations [83].
Data triangulation is encouraged but not mandatory in case study research. As Stake notes, combining multiple data sources strengthens the substantiation of constructs and hypotheses, yet practical constraints necessarily shape feasible designs [82]. We encountered such constraints.
4.5 Data Analysis
After transcription and anonymization, we began with data familiarization [90]. To analyze the data, first, we used an interpretive [27] and exploratory approach for initial data analysis, as recommended for studies characterized by uncertainty [62]. Second, because our work is a usability study, we situated the exploratory findings in a usability context by applying a formative usability framework retrospectively to our inductively generated results. This retrospective mapping of emergent themes onto existing constructs is an established interpretive strategy in qualitative analysis. This method uses prior constructs to structure and extend inductively derived findings while preserving empirical grounding (e.g., [30, 33, 18]).
Memoing & Codebook Creation
@@ 4.5 Data Analysis
We began with the initial 17 transcripts. Two analysts independently memoed a subset of six anonymized transcripts across three rounds of review and discussion, following the recommendations of O’Connor & Joffe [59] and Campbell et al. [40]. This process improved the reliability of the analysis while making efficient use of data resources. [^2] Although there is no consensus on the ideal proportion of data to use, O’Connor and Joffe note that selecting 10–25% is typical [59]. We randomly selected six transcripts ( $\sim$35% of our dataset) to develop our initial codebook. After each round, we discussed emerging codes in meetings, which produced 506 coded memos in total. We reviewed these memos to obtain a codebook of 36 categories.
@@
To assess the reliability of the coding process, we combined the techniques outlined by Maxam & Davis [54] and Campbell et al. [40]. Specifically, we randomly selected an unanalyzed transcript, which we coded independently. Following Feng et al.’s [29] recommendation, we used the percentage agreement measure to evaluate consistency, as both analysts had contributed to developing the codebook. We obtained an 89% agreement score, suggesting a stable coding scheme [40, 59]. We resolved disagreements by discussion.
Codebook Revision
@@
Following the reliability assessment of our codebook and given the substantive agreement recorded, one of the analysts proceeded to code the remaining transcripts in two batches of six and five transcripts per round, respectively. In the first batch, six new code categories were introduced, which were subsequently reviewed by the other analyst. The same analytical process was applied to the final six transcripts, leading to the addition of four new code categories, which were subsequently reviewed by the other analyst. Notably, of the 10 codes added during this phase of codebook refinement, only three provided new insights into the data. By mutual agreement between both analysts, these were categorized as minor codes.
@@
To assess the adequacy of our sample size and analysis, we calculated code saturation for each category in our codebook. Following the recommendations of Guest et al. [32], we measured the cumulative number of new codes introduced in each interview (Figure˜8). We saw saturation after interview #12.
Thematic Analysis & Usability Analysis
@@
Thematic Analysis: Each interview yielded a median of 15 codes. Next, following Braun & Clarke [13], we derived initial themes from our code categories (see Table˜1 for protocol focus). This reflexive thematic analysis provided an inductive account of participants’ experiences and practices.
Usability Framework Analysis: From the emergent themes, we interpreted our findings using a formative usability framework. Lacking formative usability studies set in similar contexts in the Software Engineering literature, we followed advice [52, 31, 79] to seek theory from adjacent disciplines.
Surveying candidate frameworks beyond SE, we found three groups: formative usability assessments of software systems [86]; applications of technology systems in other disciplines [47, 95, 41]; and general usability frameworks [89]. Among these, we selected Cresswell et al.’s four-factor TPOM framework (Technology, People, Organization, Macro-environment[41] because it best aligned with the shape of our inductively generated themes (see Figure˜1). Our intent was to synthesize those themes into higher-level conceptual groupings that correspond to established usability determinants of adoption. The TPOM categories also suggest interventions, e.g., People issues may not resolve via Technology.
We then conducted a framework-informed analysis, mapping the emergent themes to the TPOM factors to refine and structure our interpretations.
5 Results
@@
We begin by presenting a Sankey diagram (Figure˜3) that illustrates the distribution of our participants between Sigstore and non-Sigstore users, and how their teams changed signing tool usage over time. Many participants (6) transitioned from using no tools to adopting Sigstore. Others transitioned from existing tools to Sigstore: 3 from implementations that had some combinations of GPG/PGP (of which one practitioner reported combining PGP with Skopeo[61]), 2 from Notary, and 2 from proprietary or internal tools. 4 subjects did not adopt Sigstore.
@@
To understand the usability concerns that Sigstore adopters face; and why practitioners did (and did not) change tools, we structure the results by RQ using Cresswell’s framework.
5.1 RQ1: Strengths & Weaknesses of Sigstore
@@
5.1.1 Sigstore’s Strength
We extracted the benefits our participants experienced while using Sigstore from their responses. These reflections capture their firsthand experiences with Sigstore. We grouped their responses into eight categories as summarized in Table˜3. When mapped to the Cresswell framework, these factors fell into two categories: Technological and Macroenvironmental.
Table 3: Summary of Practitioner-Reported Advantages of Using Sigstore. Practitioners cited Sigstore’s Technological and Macroenvironmental factors as advantages.
@@ 5.1.1 Sigstore’s Strength
Ease of Use 8 subjects
1. Signing Workflow & Verification P1, P2, P7, P9, P14-17
2. Setting up with automated CI/CD actions P9
3. No key distribution problems P1
Use of Short-lived Keys & Certificate 3 subjects P2, P3, P15
Signer ID Management 4 subjects
1. Use of OIDC(Keyless) to authenticate signers P3, P5, P10, P12
Compatibility with Several New Technologies 4 subjects
1. Integrability with SLSA build P16
2. Integrability with several container registries/technologies P12, P14
3. Integrability with several cloud-native applications P15
Precence of a Transparency Log 3 subjects
1. Transparency logs increase security P5, P14
2. Evaluation of signing adoption using logs P9
Bundling Signatures With Provenance Attestations 2 subjects P3, P4
Reliability of Service 1 subject P7
@@ 2. Evaluation of signing adoption using logs P9
Free/Open-Source 2 subjects P7, P17
@@
We identified five factors:
Ease of use: This was the most frequently discussed advantage of Sigstore over other tools. Participants commonly highlighted the simplicity of the Sigstore workflow. As P7 (security tools, cloud security) put it, “ I don’t really know of any other options that provide the same conveniences that Sigstore provides… It really is just one command to sign something and then one more command to verify it, and so much of the work is handled by Sigstore.” Other ease-of-use advantages mentioned were the seamless integration of the Sigstore workflow into CI/CD automation using GitHub Actions and the reduced complexity of key management.
@@
Compatibility with Several New Technologies: A portion of our participants reported that Sigstore’s integration and compatibility with a wide range of modern technologies provided a significant advantage. The mentioned technologies included security tooling such as SLSA build attestations, as well as containerization technologies and cloud computing applications. In most cases, the resources required to set up connections to these technologies were minimal or nonexistent. P12 (consultancy, cloud OSS security) states, “a lot of clients are getting into signing their containers… Sigstore is the easy choice because it just works on any registry, and integrating it to Kubernetes is easy.”
Use of Short-lived Keys/Certificates: Participants appreciated Sigstore’s enforcement of short-lived keys and certificates. Some signing tools require users to manually set expiration times for cryptographic elements such as keys and certificates at the time of generation. That customizability also introduces the risk of long-lived cryptographic materials, increasing the likelihood of compromise. Participants preferred Sigstore’s automated management of key and certificate lifetimes. As P2 (SAAS, SSC security tool) stated, “PGP, you still have to figure out the key distribution problem….Sigstore uses short-lived keys, that solves a lot of those issues.”
Presence of a Transparency Log: Adopters valued the ability to audit signing actions. Sigstore’s transparency log provides a tamper-resistant record of all changes to an artifact’s metadata, a functionality not offered by other signing systems. {quoting} P9 (SAAS, SSC security): “From a higher level, we’re going to use a transparency log to determine what was actually signed and what was written into that transparency log.”
@@
Free/Open-Source: Sigstore’s open-source nature was also highlighted by participants as a key advantage. They noted that this reduces both setup and maintenance barriers, particularly for users leveraging Sigstore’s public deployments. {quoting} P7 (security tools, cloud security): “I think it’s amazing that Sigstore is free and public” This enthusiasm reflects both the appeal of free/open-source (FOSS) software and the practical benefits of public accessibility. However, note that open-source $\neq$ free; enterprises may pay for private deployments, support, and integration.
5.1.2 Sigstore’s Weaknessses
We summarize the difficulties reported by participants in using Sigstore in Table˜4. Most of these weaknesses mapped to multiple usability factors.
Table 4: Summary of Practitioner-Reported Difficulties Using Sigstore. Difficulties are grouped by Cresswell factors.